Question 2 Table 8-1 Barrett, Inc., acquired a building and the 2 acres of land on which it is located. The total purchase price was $1,000,000. For valuation purposes, the company contacted three local commercial real estate agents, who gave the following valuation estimates: Land Building Donna Gilroy $ 450,000 $1,050,000 Abby Pamer $ 600,000 $ 900,000 Megan Mallony $ 300,000 $1,200,000 Referring to Table 8-1, if Barrett, Inc., used the valuation made by Megan Mallony, and assuming it paid cash for the land and building, what journal entry would Barrett, Inc. make to record the purchase? Land 400,000 Building 600,000 Cash 1,000,000 Land 300,000 Building 1,200,000 Cash 1,500,000 Land 300,000 Building 1,200,000 Cash 1,000,000 Gain on Purchase of Assets 500,000 Land 200,000 Building 800,000 Cash 1,000,000 Land 300,000 Building 700,000 Cash 1,000,000
Answer #1
Valuation by Megan Mallony:
$
Land 300000
Building 1200000
Total 1500000
Ratio of land=300000/1500000=3/15=1/5
Ratio of Building=1200000/1500000=12/15=4/5
Ratio of Land and Building=1/5:4/5=1:4
Purchase price=$1000000
Purchase price is allocated in the ratio of 1:4 between land and building
Value of land=1000000*1/5=$ 200000
Value of Building=1000000*4/5=$ 800000
Journal entry:
Debit Credit
Land 200000
Building 800000
Cash 1000000